OpenStudy (togundeji):

5. The measures of the angles of a triangle are in the extended ratio 17 colon 16 colon 12. Which is the measure of the largest angle? (1 point) 34° 45° 48° 68°

OpenStudy (togundeji):

@aveline can you help in this one?

OpenStudy (aveline):

17x+16x+12x=180 degrees 45x=180 degrees x=4 degrees Therefore, the largest angle is 17x=17*4=68 degrees

OpenStudy (togundeji):

Thank you! Can you explain a little more?

OpenStudy (togundeji):

@aveline

OpenStudy (aveline):

Yeah

OpenStudy (aveline):

We know that all triangles have a total of 180 degrees. We know that our angle measures are 17:16:12, or 17x, 16x, and 12x. To find the value of x, we set the total of all the angles to 180: 17x+16x+12x=180 45x=180 x=4 degrees Now we have the value of x, so we can find all of the angles of the triangle. 12(4)=48 degrees 16(4)=64 degrees 17(4)=68 degrees We can double check our answer: 48+64+68=180 degrees

OpenStudy (togundeji):

ohhh okay I see!

OpenStudy (togundeji):

THANK YOU! lo! l I'll try and do the next problem on my own, and if I don't get it STILL I'll ask for help.